OpenStudy (anonymous):

2. You are trying to decide how to most efficiently use your oven. You do not want the oven running at a high temperature when you are not baking, but you also do not want to waste a lot of time waiting for the oven to reach the desired baking temperature. The instruction manual on the industrial oven suggests the oven temperature will increase by 40 degrees Fahrenheit per minute. When the oven is initially turned on, the temperature is 65 degrees Fahrenheit. Write an expression (using a variable to represent the minutes) and then calculate what the temperature of the oven will

OpenStudy (lilg132):

is that question complete?

OpenStudy (anonymous):

T=65+40*t Where T=Temperature in Fahrenheits t=time in minutes. That's the expression which relates temperature and time...

OpenStudy (anonymous):

i believe i got it..but do correct me if I am wrong: 40x + 5 = y; x represents min, and y is overall temperature of the oven after 9 minutes. so this is what i got x=9, then, multiplying 40x9+65 and y=425

OpenStudy (lilg132):

yes you are right :)

OpenStudy (lilg132):

40 x + 65 = y i think you made a mistake you wrote 40x + 5 = y

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

so how would i write that expression...

OpenStudy (anonymous):

oh wait..sorry, that is what i meant 40x+65=y

OpenStudy (lilg132):

yes you are right you just made a mistake when you were typing nothing to worry about i can see you understood correctly

OpenStudy (anonymous):

my number 6 button was sticking..

OpenStudy (anonymous):

yesssss......i am starting to get it...although, i got one more to do and not sure how to do this one... 5. When managing a business, it is important to take inventory of where your money is spent. You have a monthly budget of $5,500. Refer to the table below and answer the questions that follow. Round your answers to the nearest tenth of a percent. Category Cost Percentage Labor 39% Materials 20% Rent and utilities $1,835 Miscellaneous $420 Total $5,500 100% a. What percentage of the total monthly budget is spent on rent/utilities? b. What percentage of the total monthly budget is spent on miscellaneous items? c. How much do materials cost monthly? d. How much does labor cost monthly?

OpenStudy (anonymous):

this is what i came up with; gonna give my answers labor 2145 materials 1100 rent and utilities 3% misc 13%

OpenStudy (lilg132):

ok so to work out question a) we have to find out what 1 percent is what you need to do is get the total amount which is $5,500 which is the same as 100% now to work out the percentage of $1,835 we need to divide the amount spent on rent and utilities by the total amount so we will do this: $1,835 / $5,500 = 0.3336363636 but to get that as a percentage you will need to multiply that by 100 so it will be like this : 0.3336363636 x 100 = 33.36363636 % but it says round to nearest tenth of a percent so it should be 33.3%

OpenStudy (lilg132):

tell me if you understand how i did that

OpenStudy (anonymous):

Oh ok, yes, i understood, i didn't multiply that by 100. thats where i did a mistake..

OpenStudy (anonymous):

so for the miscelaneous i need a percentage. I need to divide 420/5500=0.076, round to the nearest tenth, would that be 10 percent?

OpenStudy (anonymous):

8% sorry

OpenStudy (lilg132):

no multiply by 100 and every number after the decimal points is tenths hundreds so e.g 0.5 is tenth 0.55 is hundreds 0.555 is thousands

OpenStudy (anonymous):

i think i got it now...i rounded off the wrong one... so 0.078 would be 8%, now my percentage all equals to 100% and all my cost totals to $5500

OpenStudy (anonymous):

0.076*

OpenStudy (lilg132):

ok so to work out question b) what you need to do is get the total amount which is $5,500 which is the same as 100% now to work out the percentage of $420 we need to divide the amount spent on miscellaneous by the total amount so we will do this: $420 / $5,500 = 0.763636363636 but to get that as a percentage you will need to multiply that by 100 so it will be like this : 0.763636363636 x 100 = 7.6% but it says round to nearest tenth of a percent so it should be 7.6%

OpenStudy (anonymous):

got it..my rounding off is totally wrong..ugh..

OpenStudy (lilg132):

no but i think you are correct in what you said so if you round up 7.6 it will be 8 33.3 rounded down should be 33 that will all add up to 100% :)

OpenStudy (lilg132):

otherwise my way would equal to 99.9 % so that must be wrong

OpenStudy (anonymous):

oh ok...so there will be no decimals..got it..

OpenStudy (anonymous):

another medal for you...woohoo. i got another question, but should i start another question? i did my worksheet, but i want to make sure i did it right and if i wrote my answers correctly in algebra expessions

OpenStudy (lilg132):

yes start another page

OpenStudy (anonymous):

i did get 100%, thats what my total came out too

OpenStudy (anonymous):

ok